The length of a rectangular tennis court is five feet more than twice the width. The perimiter is 82 feet. Write an equation, and solve the equation to find the width.
Answer by checkley77(12844)

L=2W+5
2L+2W=82
2(2W+5)+2W=82
4W+10+2W=82
6W=82-10
6W=72
W=72/6
W=12 IS THE WIDTH.
L=2*12+5
L=24=5
L=29
PROOF:
2*29+2*12=82
58+24=82
82=82